You have won the lottery at the beginning o f this year.  You have the option of receiving a lump sum payment (now) of $5,000 or receiving the cash over three years respectively: Year 1 $2,000; Year 2 $2,500; Year $3,000. You currently receive a risk-free rate of return of 3% in your high yield savings bank account.  When considering the time value of money, what option is more valuable.   Hint: Create a cash flow schedule for the yearly payments and compare against the future value of the lump sum payment received today.
Answer #1

We need to compare the future value of following cash flows and choose the higher one.

Year Cash flow 1 Cash flow 2
0 5000 0
1 0 2000
2 0 2500
3 0 3000

Future value of Cash flow 1 at the end of year 3 = PV*(1+r)^3

FV1 = 5000*1.03^3 = $5463.635

FV2 = 2000*1.03^2 + 2500*1.03 + 3000 = $7696.8

Since, option 2 have higher future value, this option is more valuable.